Update: This was fixed for Stack Overflow yesterday (all other sites were already updating correctly). There was an additional problem with Stack Overflow auth that took some time to track down, ultimately stemming from the cascading failure of auth due to ephemeral port opening failures on the firewalls between that SQL server and the domain controllers earlier in the week. That left the auth sessions in a very weird state but also led us to other issues, such as the SPNs not creating correctly and auth failing back to NTLM requiring the pass-through hop rather than Kerberos that should have been in play (and now is). We took the afternoon and fixed every issue we could find to ensure smooth operation from here on out. Apologies that took a while, this wasn't a quick fix and far from the only thing on the list we're working on while in maintenance mode for the other data center.
There was an error authing to the primary servers to pull down the updated data, which was just really bad timing with another update/reboot. A fresh import is running now and should take a few hours.